The sum of Carl's and Patricia's ages is 30. Thirteen years ago, Patricia was three times as old as Carl. How old is each now.
This is one of the many questions on my homework assignment. I have no clue how to figure out these problems. Please help me learn to set it up correctly.

Answer by awang1996(50)

The sum of Carl's and Patricia's ages is 30. Thirteen years ago, Patricia was three times as old as Carl. How old is each now.


c+p=30
3(c-13)=p-13
3c-39=p-13
3c-26=p
c+3c-26=30
4c-26=30
4c=56
c=14
p=16

Carl:14
Patricia:16